# Hyperparameter Optimization for Randomized Algorithms: A Case Study for Random Features

### v2

@@ 1 Introduction
Nonconvex optimization is ubiquitous in machine learning. In deep learning, it most frequently arises when training a neural network. This involves adjusting the weights and biases of the network to minimize a cost function. Alternatively, one can greatly simplify this optimization task by replacing the inner network weights with fixed samples from a probability distribution (Huang et al., 2006; Scardapane and Wang, 2017) and only learning the neural network’s linear output layer. With a suitable choice of loss function, such as a (possibly regularized) least squares functional, this optimization problem becomes a convex program. In the least squares example, its explicit solution involves the inversion of a linear system. Such an approach reduces approximation power, but does not have to deal with the challenges that come with nonconvexity. A large body of empirical evidence demonstrates that these convex reductions display remarkable performance in streaming large data sets at moderate processing and memory costs during training, partly because they do not require backpropagation (Chattopadhyay et al., 2020; Le et al., 2013; Sun et al., 2015; Bollt, 2021). Performance on (often high-dimensional) problems with little data available for training is less explored (Gauthier et al., 2021; Lanthaler and Nelsen, 2023).
@@
Naturally, there is a hidden price attached to such randomized approximations: one must pose additional structure on the problem in the form of the probability distribution $\mu$ from which the neural network weights are sampled. Critically, approximation accuracy is strongly tied to the choice of $\mu$. Unsystematic choice of $\mu$ may misrepresent performance or decrease robustness and usability. Indeed, practitioners often resort to unsystematic approaches such as manual tuning, grid searches, or fixed samples that lead to user bias, scalability issues, and overfitting, respectively. One systematic method that the present paper advocates for is to (i) choose a $u$-parametrized family $\{\mu_{u}\}_{u\in\mathcal{U}}$ of probability distributions, (ii) design a regularized stochastic optimization problem for the hyperparameter $u$, and (iii) optimize over $u\in\mathcal{U}$ to achieve consistent performance over different random realizations of the algorithm. Stage (iii) will be referred to as training for the remainder of this paper. [^1] We will also write hyperparameter training for stage (iii) to distinguish it from traditional training of machine learning models with hyperparameters held fixed. The simplicity of the convex optimization of final layer weights is now contrasted by a nonconvex stochastic optimization task for the optimal hyperparameters.
@@
Motivated by scientific and engineering contexts, in this paper we instantiate the preceding systematic optimization pipeline for the specific machine learning task of function emulation from a small noisy data set. We apply the framework to one randomized algorithm, the method of random features (RFs, Rahimi and Recht, 2007). We benchmark the RF approach with the current state-of-the-art tool for such tasks, the kernel-based method of Gaussian processes (GPs, Williams and Rasmussen, 2006). The underlying hyperparameter training problem for the RF method involves optimally adapting the RF sampling distribution $\mu_{u}$ to data. This is equivalent to learning a data-adapted kernel or Bayesian prior distribution. Indeed, RFs, which in special cases can be viewed as neural networks with randomly sampled weights (e.g., “extreme learning machines” from Huang et al. (2006)), by construction can also be viewed as low-rank approximations to GPs. A broad chronology of the developments and performance gains from RF approximation of kernel and GP methods can be found in the work of Liu et al. (2022).
@@
The majority of existing studies do not systematically optimize the RF sampling distribution $\mu_{u}$. Of those that do, the approach of Yang et al. (2015) is closely related to our own. The authors adopt a GP perspective on RFs by deriving a cost function from empirical Bayes arguments. The work fixes the number and specific realizations of the sampled features. Although this produces good results, their approach only fits to the individual sample and not its law, which prohibits future resampling and flexible computational cost adjustments. Other fixed-feature approaches use random projections (Hamid et al., 2014) or discrete kernel alignment (Sinha and Duchi, 2016) to create an optimal weighting of a pre-generated set of RFs. More modern methods learn a generative model for the feature distribution (Li et al., 2019; Falk et al., 2022); this results in more flexible RF methods. Another line of research derives explicit formulas for the “optimal” RF distribution based on minimization of upper bounds on the approximation error (Kammonen et al., 2020, 2023). The authors then approximately sample from this inaccessible distribution with Markov chain Monte Carlo (MCMC) methods. To summarize, most existing works that do adopt systematic hyperparameter calibration pipelines must exploit special properties and structure of the distribution family $\{\mu_{u}\}$. In contrast, the present paper introduces a widely applicable hyperparameter learning methodology that treats $\{\mu_{u}\}$ in a black-box manner.
@@
The optimization workhorse of this paper is the ensemble Kalman inversion (EKI) algorithm (Iglesias et al., 2013; Schillings and Stuart, 2017; Calvello et al., 2025), one of a family of derivative-free algorithms designed to solve nonlinear optimization problems that arise from Bayesian inversion. It can be configured to solve for both a maximum likelihood estimator or a maximum a posteriori estimator. For linear inversion with Gaussian priors and Gaussian observational noise models, the EKI algorithm converges to the true point estimator of the Bayesian posterior distribution. As a particle-based method, EKI can be derived from the ensemble Kalman filtering (EnKF) literature (Chen and Oliver, 2012; Iglesias et al., 2013) and, along with its variants, has been used successfully in a wide range of optimization and inversion tasks (Böttcher, 2023; Cleary et al., 2021; Dunbar et al., 2021, 2022a, 2022b; Huang et al., 2022; Kovachki and Stuart, 2019; Xiao et al., 2016). In part, this success is due to the method’s robustness to rough optimization landscapes and the inheritance of modifiers (e.g., localization, sampling error correction, covariance inflation) and variants (e.g., square root inversion, unscented inversion, sparsity-promoting inversion) from the decades-long history of EnKF in meteorological applications. Since EKI treats the hyperparameter-to-output map as a black-box (in particular, without access to derivatives of the map), it allows for easy prototyping and modularity over different families of features, hyperparameters, and loss functions.
@@ 1.1 Contributions
    This is in contrast to much of the broader machine learning literature, which is more concerned with scaling performance on massive data sets (Liu et al., 2022) and more reliant on approaches such as exact diagonalization to tackle multi-output problems.
@@
    We test the new machine learning tools on three scientifically-minded benchmark problems. The trained tools are used to emulate scalar-valued functions for global sensitivity analysis, integrate the state a chaotic Lorenz 63 dynamical system, and accelerate a Bayesian parameter estimation problem from the atmospheric sciences by emulating the underlying forward map. In all three applications, we compare our approach with a current gold-standard GP-based methodology.
@@
    We document and actively maintain all tools and examples in open-source, registered, Julia-language Github repositories (Dunbar et al., 2022b, 2024).
@@ 1.2 Outline
The remainder of this paper is organized as follows. Section 2 introduces Bayesian regression with RFs from a GP perspective for both scalar-valued and vector-valued function learning settings. Section 3 motivates the RF hyperparameter learning problem from empirical Bayes ideas. It also presents the black-box EKI optimizer that is championed in this work and a random objective function that is amenable to EKI. Section 4 applies the trained RF emulators to three scientific applications. The paper concludes with a final discussion in Section 5. Appendix A contains the additional details regarding the numerical experiments from Section 4.
@@ 2.1.2 Random Feature Regression
Several scalable approximations to GPR have been proposed in the literature. These include, but are not limited to, sparse variational approaches based on inducing points (Titsias, 2009), Nyström subsampling (Sun et al., 2015), and RFs (Rahimi and Recht, 2007). We focus on RFs and in particular the method of RFR. We interpret RFR as an approximation to GPR obtained by first replacing the GP prior covariance $K$ with a low-rank approximation $K_{M}$ and then performing exact GPR inference (5) with this low-rank covariance function. We now describe this procedure in detail.
@@ 2.2.1 Vector-Valued Gaussian Process Regression
The factor $K(x,X)=K(X,x)^{\top}\in(\mathbb{R}^{p\times p})^{1\times N}$ is interpreted as a “row vector” with matrix entries. The concatenated data $Y=\{y_{n}\}_{n=1}^{N}$ and prior mean entries $\overline{F}(X)$ are viewed as length $N$ vectors with each entry taking value in $\mathbb{R}^{p}$. The formulas (20) follow by whitening (19) (i.e., pre-multiplying by $\Sigma^{-1/2}$) and either applying a change of variables argument to an existing white noise result (Owhadi, 2022, Theorem 5.3) or applying a different result (Knapik et al., 2011, Proposition 3.1) alongside the Doob–Dynkin lemma.
@@ 2.2.2 Vector-Valued Random Feature Regression
where $\ell\in[M]$. The $\Sigma$-weighted system Gram matrices in (25) are still only of size $M$ by $M$. These are much cheaper to invert than the vector-valued GPR system matrices whenever $M\ll Np$; this is often the case in practice. Table 1 summarizes the computational tradeoffs. We observe that the complexity of vector-valued GPR depends directly on $Np$, with cubic scaling in $Np$ offline and quadratic scaling online for storage. On the other hand, the complexity of RFR depends instead on $M$ instead of $Np$, with similar scaling in this variable. Thus, RFR is computationally advantageous whenever $M\ll Np$, while still maintaining the same average prediction accuracy as vector-valued GPR (Lanthaler and Nelsen, 2023).
@@ 3 Hyperparameter Learning for Random Feature Regression
Recall that in the development of Section 2, the RF pair $(\varphi,\mu)$ is assumed given. An often overlooked practical consideration of RFR—or other algorithms based on RFs—is how to choose the pair $(\varphi,\mu)$. This is the natural analog to choosing the covariance kernel $K$ in GPR. To address this kernel selection problem, we begin in Subsection 3.1 by adapting the popular empirical Bayes approach for learning priors in GPR to the stochastic setting of RFR. This approach comes with several unique challenges that we address with the derivative-free optimization method known as EKI. We describe the EKI algorithm in Subsection 3.2. In Subsection 3.3, we cast the covariance kernel learning problem as a statistical inverse problem so that it may be solved with EKI. Subsection 3.4 instantiates the methodology with an expressive family of parametrized distributions $\{\mu_{u}\}$ that also performs well in practice.
@@ 3.1 Empirical Bayes Motivation
$G_{M}^{(u)}\coloneqq\biggl{[}\frac{1}{M}\sum_{n=1}^{N}\varphi\bigl{(}x_{n};\theta_{\ell}^{(u)}\bigr{)}^{\top}\Sigma^{-1}\varphi\bigl{(}x_{n};\theta_{m}^{(u)}\bigr{)}+\delta_{\ell m}\biggr{]}_{\ell\in[M],m\in[M]},$ (29)
@@
The calculations in Subsection 2.2.2 show that the RFR posterior mean coefficients $\beta=\beta^{(u)}\in\mathbb{R}^{M}$ from (25) satisfy $\beta_{m}^{(u)}=\sum_{n=1}^{N}\varphi(x_{n};\theta_{m}^{(u)})^{\top}\alpha_{n}^{(u)}$. Using this and the fact that $B_{\Sigma}\alpha^{(u)}=Y-K_{M}^{(u)}(X,X)\alpha^{(u)}$, we define our practically computable random empirical Bayes objective function to be
${L}^{(\mathrm{EB})}_{M}(u)\coloneqq\frac{1}{M}\big{\lVert}\beta^{(u)}\big{\rVert}^{2}_{\mathbb{R}^{M}}+\Big{\lVert}B_{\Sigma}^{-1/2}\bigl{(}Y-\overline{F}^{(N)}_{M}(X;u)\bigr{)}\Big{\rVert}^{2}_{(\mathbb{R}^{p})^{N}}+\log\det\bigl{(}G_{M}^{(u)}+I_{M}\bigr{)}.$ (34)
This expression is equal to $\widetilde{L}^{(\mathrm{EB})}_{M}$ but with the term $N\log\det(\Sigma)$ subtracted off because it is independent of $u$. The map $\overline{F}^{(N)}_{M}({\,\cdot\,};u)$ in (34) is the RF posterior mean corresponding to feature pair $(\varphi,\mu_{u})$.
@@ 3.2 Ensemble Kalman Inversion
EKI (Iglesias et al., 2013; Schillings and Stuart, 2017; Calvello et al., 2025)—and its variants (Huang et al., 2022; Garbuno-Inigo et al., 2020a, b)—is a family of tools for stochastic optimization. More explicitly, EKI is a method for solving Bayesian inverse problems. In the following exposition, we first describe the basics of the algorithm to solve such problems and then cast a new inverse problem whose solution is related to minimization of (34).
The underlying inverse problem is as follows. Assume a system produces observable output under a parameter-to-observable map $\mathcal{G}$ that is random with Gaussian marginals. The map can be written as $\mathcal{G}\colon u\mapsto\mathcal{G}(u)$ such that $\mathcal{G}(u)\sim\mathsf{N}(\overline{\mathcal{G}}(u),\Gamma(u))$ for some (possibly $u$-dependent) covariance matrix $\Gamma=\Gamma(u)$. Here the “overbar” notation represents the mean of the distribution of $\mathcal{G}(u)$. The relationship between the parameters and observable $z$ can be written as
$z$ $=\mathcal{G}(u),\quad\text{where}\quad\mathcal{G}(u)\sim\mathsf{N}\bigl{(}\overline{\mathcal{G}}(u),\Gamma(u)\bigr{)},$ (35)
that is,
$z$ $=\overline{\mathcal{G}}(u)+\eta,\quad\text{where}\quad\eta\sim\mathsf{N}(0,\Gamma(u)).$ (36)
The form (36) emphasizes that the relationship can be written in terms of a deterministic map plus additive noise. By endowing the input $u\in\mathcal{U}$ with a Bayesian prior distribution, and given a particular observation $z=z^{\star}$, we can informally state two optimization tasks (i.e., inverse problems): find an optimal $u^{\star}=u^{(\mathrm{MLE})}$ or $u^{\star}=u^{(\mathrm{MAP})}$, where
@@
Here $P(w)$ denotes the probability density function of a finite-dimensional random variable $w$. In (37), $P(z^{\star}\,|\,u)$ is the data likelihood for the observation $z^{\star}$ arising from any input $u$ and $P(u)$ represents the prior on $u$. Commonly, optimization frameworks evaluate (36) to compute $P(z^{\star}\,|\,u)$ and so require approximation of $\overline{\mathcal{G}}(u)$, a computationally demanding task (e.g., via sampling). On the other hand, EKI solves for $u^{\star}$ from (35) and requires only evaluations of $u\mapsto\mathcal{G}(u)$.
The basic EKI algorithm makes another modeling assumption typical of the ensemble Kalman literature by insisting that the prior is a Gaussian $\mathsf{N}(m,C)$. Under this condition, it holds that
@@
where the norms are induced by the Euclidean inner product. The algorithm represents the Gaussian distributions empirically from samples (ensemble members) and evolves the ensemble with a Kalman-type update rule. First, the critical EKI hyperparameter is chosen, the ensemble size $J$. The initial ensemble can be written for each member $j\in[J]$ as $u^{(j)}(t_{0})\sim\mathsf{N}(m,C)$. Given a sequence $\{t_{n}\}$ of artificial time values, the update rule for EKI at iteration $n\in[n_{0}-1]$ is given by
@@
The time step $\Delta t_{n+1}$ is usually constant, e.g., $\Delta t_{n+1}\equiv\Delta t=1$, and $u^{(j)}(t_{n})$ is often also denoted by $u^{(j)}_{n}$. Writing $\overline{u(t)}\coloneqq\frac{1}{J}\sum_{j=1}^{J}u^{(j)}(t)$ and similarly [^6] Note the difference in notation: $\overline{\mathcal{G}}(u)$, the mean of the distribution at $u$, versus $\overline{\mathcal{G}(u)}$, the finite ensemble average of evaluations of $\mathcal{G}$ over $u_{j}$ for $j\in[J]$. for $\overline{\mathcal{G}(u)}$, for any $t$ the empirical covariance matrices in (39) are given by
@@
In the setting of a linear parameter-to-observable map $\mathcal{G}$, iterating until time $t_{n_{0}}=T=1$ will approximate the posterior distribution of the Bayesian inverse problem (Calvello et al., 2025), that is, the mean and covariance of the final ensemble will accurately represent the posterior mean and covariance. In the nonlinear case, iterating until $T=1$ (just one step if $\Delta t\equiv 1$) will approximate $u^{(\mathrm{MAP})}$. In contrast, upon sending $t\to\infty$, the dynamic will instead approach $u^{(\mathrm{MLE})}$—up to constraints imposed by the span of the initial ensemble. The ensemble is known to greatly underestimate posterior covariances (even at $T=1$), and so these will be disregarded in this investigation. Other variants of the ensemble Kalman methodology have been developed to overcome this limitation (Garbuno-Inigo et al., 2020a, b).
The algorithm (39) is the basic form of EKI. As well as aforementioned variants, there are several features that can be added for additional computational advantages. Many of these may be found in the open-source software package of Dunbar et al. (2022b) and are used later on in the numerical experiments of the present paper. Detailed EKI configurations are provided for each experiment in the appendix.

### v4

# Hyperparameter Optimization for Randomized Algorithms: A Case Study on Random Features

@@ 1 Introduction
Nonconvex optimization is ubiquitous in machine learning. In deep learning, it most frequently arises when training a neural network. This involves adjusting the weights and biases of the network to minimize a loss function. Alternatively, one can greatly simplify this optimization task by replacing the inner network weights with fixed samples from a probability distribution (Huang et al., 2006; Scardapane and Wang, 2017) and only learning the neural network’s linear output layer. With a suitable choice of loss function, such as a (possibly regularized) least squares functional, this optimization problem becomes a convex program. In the least squares example, its explicit solution involves the inversion of a linear system. Such an approach reduces approximation power, but does not have to deal with the challenges that come with nonconvexity. A large body of empirical evidence demonstrates that these convex reductions display remarkable performance in streaming large data sets at moderate processing and memory costs during training, partly because they do not require backpropagation (Chattopadhyay et al., 2020; Le et al., 2013; Sun et al., 2015; Bollt, 2021). Performance on (often high-dimensional) problems with little data available for training is less explored (Gauthier et al., 2021; Lanthaler and Nelsen, 2023).
@@
Naturally, there is a hidden price attached to such randomized approximations: one must pose additional structure on the problem in the form of the probability distribution $\mu$ from which the neural network weights are sampled. Critically, approximation accuracy is strongly tied to the choice of $\mu$. Unsystematic choice of $\mu$ may misrepresent performance or decrease robustness and usability. Indeed, practitioners often resort to unsystematic approaches such as manual tuning, grid searches, or fixed samples that lead to user bias, scalability issues, and overfitting, respectively. One systematic method that the present paper advocates for is to (i) choose a $u$-parametrized family $\{\mu_{u}\}_{u\in\mathcal{U}}$ of probability distributions, (ii) design a regularized stochastic optimization problem for the hyperparameter $u$, and (iii) optimize over $u\in\mathcal{U}$ to achieve consistent performance over different random realizations of the algorithm. Stage (iii) will be referred to as tuning or calibration for the remainder of this paper. The simplicity of the convex optimization of final layer weights is now contrasted by a nonconvex stochastic optimization task for the optimal hyperparameters.
@@
Motivated by scientific and engineering contexts, in this paper we instantiate the preceding systematic optimization pipeline for the specific machine learning task of function emulation from a small noisy data set. We apply the framework to one randomized algorithm, the method of random features (RFs, Rahimi and Recht, 2007). We benchmark the RF approach with the current state-of-the-art tool for such tasks, the kernel-based method of Gaussian processes (GPs, Williams and Rasmussen, 2006). The underlying hyperparameter tuning problem for the RF method involves optimally adapting the RF sampling distribution $\mu_{u}$ to data. This is equivalent to learning a data-adapted kernel or Bayesian prior distribution. Indeed, RFs, which in special cases can be viewed as neural networks with randomly sampled weights (e.g., “extreme learning machines” from Huang et al. (2006)), by construction can also be viewed as low-rank approximations to GPs. A broad chronology of the developments and performance gains from RF approximation of kernel and GP methods can be found in the work of Liu et al. (2022).
@@
The majority of existing studies do not systematically optimize the RF sampling distribution $\mu_{u}$. Of those that do, the approach of Yang et al. (2015) is closely related to our own. The authors adopt a GP perspective on RFs by deriving an objective function from empirical Bayes arguments. The work fixes the number and specific realizations of the sampled features. Although this produces accurate results, their approach only fits to the individual sample and not its law, which prohibits future resampling and flexible computational cost adjustments. Other fixed-feature approaches use random projections (Hamid et al., 2014) or discrete kernel alignment (Sinha and Duchi, 2016) to create an optimal weighting of a pre-generated set of RFs. More modern methods learn a generative model for the feature distribution (Li et al., 2019; Falk et al., 2022); this results in more flexible RF methods. Another line of research derives explicit formulas for the “optimal” RF distribution based on minimization of upper bounds on the approximation error (Kammonen et al., 2020, 2023). The authors then approximately sample from this inaccessible distribution with Markov chain Monte Carlo (MCMC) methods. To summarize, most existing works that do adopt systematic hyperparameter calibration pipelines must exploit special properties and structure of the distribution family $\{\mu_{u}\}$. In contrast, the present paper introduces a widely applicable hyperparameter learning methodology that treats $\{\mu_{u}\}$ in a black-box manner.
@@
The optimization workhorse of this paper is the ensemble Kalman inversion (EKI) algorithm (Iglesias et al., 2013; Schillings and Stuart, 2017; Calvello et al., 2025), one of a family of derivative-free algorithms designed to solve nonlinear optimization problems that arise from Bayesian inversion. It can be configured to solve for both a maximum likelihood estimator or a maximum a posteriori estimator. For linear inversion with Gaussian priors and Gaussian observational noise models, the EKI algorithm converges to the true point estimator of the Bayesian posterior distribution. As a particle-based method, EKI can be derived from the ensemble Kalman filtering (EnKF) literature (Chen and Oliver, 2012; Iglesias et al., 2013) and, along with its variants, has been used successfully in a wide range of optimization and inversion tasks (Böttcher, 2023; Cleary et al., 2021; Dunbar et al., 2021, 2022a, 2022b; Huang et al., 2022; Kovachki and Stuart, 2019; Xiao et al., 2016). In part, this success is due to the method’s robustness to rough optimization landscapes and the inheritance of modifiers (e.g., localization, sampling error correction, covariance inflation) and variants (e.g., square root inversion, unscented inversion, sparsity-promoting inversion) from the decades-long history of EnKF in meteorological applications. Since EKI treats the hyperparameter-to-output map as a black-box (in particular, without access to derivatives of the map), it allows for easy prototyping and modularity over different families of features, hyperparameters, and objective functions. On the other hand, algorithms that make use of higher-order derivative information tend to converge faster than those that do not. However, the derivative-free EKI algorithm is still known to approximate a gradient descent dynamic and hence acceleration methods such as momentum are readily applicable to EKI (Kovachki and Stuart, 2019, Section 4).
@@ 1.1 Contributions
    This is in contrast to much of the broader machine learning literature, which is more concerned with scaling performance on massive data sets and more reliant on approaches such as exact diagonalization to tackle multi-output problems.
@@
    We test the new machine learning tools on three scientifically-minded benchmark problems. The tools are used to emulate scalar-valued functions for global sensitivity analysis, integrate the state a chaotic Lorenz 63 dynamical system, and accelerate a Bayesian parameter estimation problem from the atmospheric sciences by emulating the underlying forward map. In all three applications, we compare our approach with a current gold-standard GP-based methodology.
@@
    We document and actively maintain all tools and examples in open-source, registered, Julia-language GitHub repositories (Dunbar et al., 2022b, 2024) with documentation at:
@@ 1.2 Outline
The remainder of this paper is organized as follows. Section 2 introduces Bayesian regression with RFs from a GP perspective for both scalar-valued and vector-valued function learning settings. Section 3 motivates the RF hyperparameter learning problem from empirical Bayes ideas. It also presents the black-box EKI optimizer that is championed in this work and a random objective function that is amenable to EKI. Section 4 applies the tuned RF emulators to three scientific applications. The paper concludes with a final discussion in Section 5. Appendix A contains additional details regarding the numerical experiments from Section 4.
@@ 2.1.2 Random Feature Regression
Several scalable approximations to GPR have been proposed in the literature. These include, but are not limited to, sparse variational approaches based on inducing points (Hensman et al., 2015, 2018; Titsias, 2009), Nyström subsampling (Sun et al., 2015), and RFs (Rahimi and Recht, 2007). We focus on RFs and in particular the method of RFR. We interpret RFR as an approximation to GPR obtained by first replacing the GP prior covariance $K$ with a low-rank approximation $K_{M}$ and then performing exact GPR inference (5) with this low-rank covariance function. We now describe this procedure in detail.
@@ 2.2.1 Vector-Valued Gaussian Process Regression
The factor $K(x,X)=K(X,x)^{\top}\in(\mathbb{R}^{p\times p})^{1\times N}$ is interpreted as a “row vector” with matrix entries. The concatenated data $Y=\{y_{n}\}_{n=1}^{N}$ and prior mean entries $\overline{F}(X)$ are viewed as length $N$ vectors with each entry taking value in $\mathbb{R}^{p}$. The formulas (20) follow from whitening (19) by pre-multiplying by $\Sigma^{-1/2}$ and then applying an existing white noise result (Owhadi, 2022, Theorem 5.3).
@@ 2.2.2 Vector-Valued Random Feature Regression
where $\ell\in[M]$ and $\delta_{\ell m}=1$ if $\ell=m$ and equals zero otherwise. The $\Sigma$-weighted system Gram matrices in (25) are still only of size $M$ by $M$. These are much cheaper to invert than the vector-valued GPR system matrices whenever $M\ll Np$; this is often the case in practice. Table 1 summarizes the computational tradeoffs. We observe that the complexity of vector-valued GPR depends directly on $Np$, with cubic scaling in $Np$ offline and quadratic scaling online for storage. On the other hand, the complexity of RFR depends instead on $M$ instead of $Np$, with similar scaling in this variable. Thus, RFR is computationally advantageous whenever $M\ll Np$, while still maintaining the same average prediction accuracy as vector-valued GPR (Lanthaler and Nelsen, 2023).
@@ 3 Hyperparameter Learning for Random Feature Regression
Recall that in the development of Section 2, the RF pair $(\varphi,\mu)$ is assumed given. An often overlooked practical consideration of RFR—or other algorithms based on RFs—is how to choose the pair $(\varphi,\mu)$. This is the natural analog to choosing the covariance kernel $K$ in GPR. To address this kernel selection problem, we begin in Subsection 3.1 by adapting the popular empirical Bayes approach for learning priors in GPR to the stochastic setting of RFR. This approach comes with several unique challenges that we address with the derivative-free optimization method known as EKI. We describe the EKI algorithm in Subsection 3.2. In Subsection 3.3, we cast the covariance kernel learning problem as a statistical inverse problem so that it may be solved with EKI. Subsection 3.4 instantiates the methodology with an expressive family of parametrized distributions $\{\mu_{u}\}$ that also performs well in practice. Subsection 3.5 acknowledges the limitations of the proposed framework.
@@ 3.1 Empirical Bayes Motivation
$G_{M}^{(u)}\coloneqq\biggl{[}\frac{1}{M}\sum_{n=1}^{N}\varphi\bigl{(}x_{n};\theta_{\ell}^{(u)}\bigr{)}^{\top}\Sigma^{-1}\varphi\bigl{(}x_{n};\theta_{m}^{(u)}\bigr{)}\biggr{]}_{\ell\in[M],m\in[M]},$ (29)
@@
The calculations in Subsection 2.2.2 show that the RFR posterior mean coefficients $\beta=\beta^{(u)}\in\mathbb{R}^{M}$ from (25) satisfy $\beta_{m}^{(u)}=\sum_{n=1}^{N}\varphi(x_{n};\theta_{m}^{(u)})^{\top}\alpha_{n}^{(u)}$. This implies that the first term on the right hand side of (33) equals $\lVert\beta^{(u)}\rVert^{2}_{\mathbb{R}^{M}}/M$. The second term may be written as $(\alpha^{(u)})^{\top}B_{\Sigma}^{\phantom{-1}}B_{\Sigma}^{-1}\,B_{\Sigma}^{\phantom{-1}}\alpha^{(u)}=\lVert B_{\Sigma}^{-1/2}(B_{\Sigma}\alpha^{(u)})\rVert^{2}_{(\mathbb{R}^{p})^{N}}$. Using the fact that $B_{\Sigma}\alpha^{(u)}=Y-K_{M}^{(u)}(X,X)\alpha^{(u)}$, we thus define our practically computable random empirical Bayes objective function to be ${L}^{(\mathrm{EB})}_{M}(u)\coloneqq\widetilde{L}^{(\mathrm{EB})}_{M}-N\log\det(\Sigma)$, which equals
${L}^{(\mathrm{EB})}_{M}(u)=\frac{1}{M}\big{\lVert}\beta^{(u)}\big{\rVert}^{2}_{\mathbb{R}^{M}}+\Big{\lVert}B_{\Sigma}^{-1/2}\bigl{(}Y-\overline{F}^{(N)}_{M}(X;u)\bigr{)}\Big{\rVert}^{2}_{(\mathbb{R}^{p})^{N}}+\log\det\bigl{(}G_{M}^{(u)}+I_{M}\bigr{)}.$ (34)
We subtracted off the term $N\log\det(\Sigma)$ (32) because it is independent of $u$. The map $\overline{F}^{(N)}_{M}({\,\cdot\,};u)=K_{M}^{(u)}({\,\cdot\,},X)\alpha^{(u)}$ in (34) is the RF posterior mean corresponding to feature pair $(\varphi,\mu_{u})$.
@@ 3.2 Ensemble Kalman Inversion
EKI (Iglesias et al., 2013; Schillings and Stuart, 2017; Calvello et al., 2025)—and its variants (Huang et al., 2022; Garbuno-Inigo et al., 2020a, b)—is a family of tools for stochastic optimization. More explicitly, EKI is a method for solving Bayesian inverse problems. In the following exposition, we first describe the basics of the algorithm to solve such problems. Then the following subsection proposes a new inverse problem whose solution is related to minimization of (34).
The underlying generic inverse problem setting is as follows. Assume a system produces observable output under a parameter-to-observable map $\mathcal{G}$ that is random with Gaussian marginals. That is, the relationship between the parameters $u$ and observable $z$ is given by
$z$ $=\mathcal{G}(u),\quad\text{where}\quad\mathcal{G}(u)\sim\mathsf{N}\bigl{(}\overline{\mathcal{G}}(u),\Gamma(u)\bigr{)}.$ (35)
In the preceding display, $\Gamma=\Gamma(u)$ is some (possibly $u$-dependent) covariance matrix, and the “overbar” notation $\overline{\mathcal{G}}$ represents the mean of the distribution of $\mathcal{G}$. Equivalently,
$z$ $=\overline{\mathcal{G}}(u)+\eta,\quad\text{where}\quad\eta\sim\mathsf{N}\bigl{(}0,\Gamma(u)\bigr{)}.$ (36)
The form (36) emphasizes that the relationship can be written in terms of a deterministic map $\overline{\mathcal{G}}$ plus additive noise. By endowing the input $u\in\mathcal{U}$ with a Bayesian prior distribution, and given a particular observation $z=z^{\star}$, we can informally state two optimization tasks (i.e., inverse problems): find an optimal $u^{\star}=u^{(\mathrm{MLE})}$ or $u^{\star}=u^{(\mathrm{MAP})}$, where
@@
Here $w\mapsto P(w)$ denotes the probability density function of a finite-dimensional random variable $w$. In (37), $P(z^{\star}\,|\,u)$ is the data likelihood for the observation $z^{\star}$ arising from any input $u$ and $P(u)$ represents the prior on $u$. Commonly, optimization frameworks evaluate (36) to compute $P(z^{\star}\,|\,u)$ and so require approximation of $\overline{\mathcal{G}}(u)$, which is a computationally demanding task in many applications (e.g., via sampling). On the other hand, EKI solves for $u^{\star}$ from (35) and requires only evaluations of the parameter-to-observable map $u\mapsto\mathcal{G}(u)$.
We now describe EKI. The basic EKI algorithm makes a modeling assumption that is typical of the ensemble Kalman literature by insisting that the prior on parameter $u$ is a Gaussian distribution $\mathsf{N}(m,C)$. Under this condition, we note that
@@
where the norms are induced by the Euclidean inner product. EKI represents the initial Gaussian distribution empirically from samples; these are called ensemble members. Let the size of the ensemble be denoted by $J$, which is a critical hyperparameter in EKI. The initial ensemble can be written for each member $j\in[J]$ as i.i.d. samples $u^{(j)}(t_{0})\sim\mathsf{N}(m,C)$. Given a sequence $\{t_{n}\}$ of artificial time values, the Kalman-type update rule for EKI at iteration $n\in[n_{\mathrm{iter}}-1]$ is given for each $j\in[J]$ by
@@
The first equation updates the parameter and the second equation represents a noisy perturbation of the observation $z^{\star}$. For further interpretations and derivations of the EKI update formulas, see Sanz-Alonso et al. (2023, Chapter 13) and references therein. The time step $\Delta t_{n+1}$ in the preceding display is usually constant, e.g., $\Delta t_{n+1}\equiv\Delta t=1$. Writing $\overline{u(t)}\coloneqq\frac{1}{J}\sum_{j=1}^{J}u^{(j)}(t)$ and similarly [^5] Note the difference in notation: $\overline{\mathcal{G}}(u)$, the mean of the distribution of $\mathcal{G}$ evaluated at $u$, versus $\overline{\mathcal{G}(u(t))}$, the finite ensemble average of evaluations of $\mathcal{G}$ over $u^{(j)}(t)$ for $j\in[J]$. for $\overline{\mathcal{G}(u(t))}$, for any $t$ the empirical covariance matrices in (39) are given by
@@
This completes the description of the basic EKI algorithm.
In the setting of a linear parameter-to-observable map $\mathcal{G}$, iterating (39) until time $t_{n_{\mathrm{iter}}}=T=1$ will approximate the posterior distribution of the Bayesian inverse problem (Calvello et al., 2025). That is, the mean and covariance of the final ensemble will accurately represent the posterior mean and covariance. In the nonlinear case, iterating until $T=1$ (just one step if $\Delta t\equiv 1$) will approximate $u^{(\mathrm{MAP})}$ in (38). In contrast, upon sending $t\to\infty$, the dynamic will instead approach $u^{(\mathrm{MLE})}$—up to constraints imposed by the span of the initial ensemble. The ensemble is known to greatly underestimate posterior covariances (even at $T=1$), and so these will be disregarded in this paper except in Section 4.3. Other variants of the ensemble Kalman methodology have been developed to overcome this limitation (Garbuno-Inigo et al., 2020a, b).
The algorithm (39) is the basic form of EKI. As well as previously mentioned variants, there are several features that can be added for additional computational advantages. Many of these may be found in the open-source software package by Dunbar et al. (2022b) and are used later on in the numerical experiments of the present paper. Detailed EKI configurations are provided for each upcoming experiment in Appendix A.